If a, b, c are in GP and the equations ax^2 + 2bx + c = 0 and dx^2 + 2ex + f = 0 have a common root,
then
d/a ,e/b ,f/c
are in

Grade:11

1 Answers

Deepak Kumar Shringi
6 years ago
If the equationsax^2+2bx+c=0and dx^2+2ex+f=0have one and only one root common
then(ae-bd)(bf-ec)=(cd-af)2use this and also replace b2=ac